OpenStudy (anonymous):

Please help me Will medal If $5700 is invested in a savings account for which interest is compounded quarterly, and if the $5700 turns into $6600 in 4 years what is the interest rate of the savings account? A 14.93% B 4.92% C 3.68% D 7.40%

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Solve for r \[\Large A = P\left(1+\frac{r}{n}\right)^{nt}\] \[\Large 6600 = 5700\left(1+\frac{r}{4}\right)^{4*4}\] \[\Large 6600 = 5700\left(1+\frac{r}{4}\right)^{16}\] \[\Large \frac{6600}{5700} = \left(1+\frac{r}{4}\right)^{16}\] \[\Large \frac{22}{19} = \left(1+\frac{r}{4}\right)^{16}\] \[\Large \left(\frac{22}{19}\right)^{1/16} = 1+\frac{r}{4}\] I'll let you finish up

OpenStudy (anonymous):

3.68% right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np